### v1

# Numerical analysis of heat transport in classical one-dimensional systems

Abstract
Numerical studies of some unidimensional systems provide evidence of finite thermal conductivity, where theory predicts a divergence with the system size. Some models are here reviewed under the working hypothesis that the energy flux across a nonequilibrium stationary state may be the sum of two contributions: the former responsible for normal conductivity, the latter accounting for the anomalous component. I conclude that in the thermodynamic limit, the diverging component is indeed present, although the crossover to a regime dominated by the anomalous component may occur at extremely long system sizes.
Finally, I study a variant of the ding-a-ling model, previously claimed to satisfy Fourier law, showing that it too exhibits a diverging conductivity in the thermodynamic limit.
1 Introduction
Since numerical evidence of anomalous heat conductivity was first given in chains of classical nonlinear oscillators [1], much progress has been made both by simulating nonequilibrium steady states and by studying fluctuations of equilibrium regimes [2, 3, 4]. On the theoretical side, the key cornerstones are based on mode coupling theories [5, 6], kinetic approaches [7, 8, 9], and fluctuating-hydrodynamic equations [10].
The typical setup where heat conduction is investigated consists of a chain of oscillators interacting via nearest-neighbour forces (mediated by a potential $V(u)$, where $u$ is the mutual distance), which exchange energy with two thermal reservoirs located at its extrema and operating at different temperatures. It is well established that whenever energy is the only conserved quantity, normal diffusion and normal conductivity are expected: i.e. the energy flux is inversely proportional to the system size $L$. A chain of oscillators interacting with an external substrate is a typical example thereof.
The presence of a second conservation law, namely linear momentum (a consequence of translational invariance), may induce a diverging, anomalous, heat conductivity. However, this divergence can arise only if a further conservation law is present: the so called total stretch, defined as the sum of the mutual particle separations $u$. In practice, this is automatically ensured when $u$ is a physical length (either longitudianal or transversal). Normal conductivity is, instead, observed when $u$ is an angular variable, as for instance in coupled rotors [11, 12]. In this case, the total stretch being the sum of all mutual angle separations can diffuse, since a variation of $2\pi$ does not have any physical implication on the dynamics. The overall scenario is well captured by the fluctuating-hydrodynamics approach [10] based on three coupled fields, associated to three conservation laws (energy, linear momentum, and total stretch).
According to theory, thermal conductivity generically diverges as $L^{1/3}$. The major exception is represented by systems with symmetric potential (and zero pressure); the most renown example is the FPUT-$\beta$ model, where $V(u)=u^{2}/2+u^{4}/4$. In such a subclass, fluctuating hydrodynamics predicts an $L^{1/2}$ divergence, while kinetic theories predict an exponent 2/5 [7, 13]. Numerical simulations confirm the latter prediction [14, 15].
Finally, there exists the special class of integrable systems; they are typically characterized by a ballistic dynamics, where the flux is independent of the system size, the reason being that the independent modes travel undamped from one to the other boundary of the chain. However, integrable systems with symmetric potentials may again display an unexpected behavior. This is, for instance, the case of the hard-point chain, where the potential $V(u)$ is an infinitely high box of width $w$. If the particles have all the same mass, each collision, occurring either when $u=0$, or $u=w$ amounts to an exchange of velocities. Hence, the dynamics is characterized by conserved quantities: the “velocitons” which jump from one to another particle, when a collision occurs. If the average interparticle distance is set equal to $w/2$, the velocitons do not move in a preferential direction; they diffuse in an anomalous way, leading to a non ballistic divergence of the heat conductivity [16].
The general scenario has been confirmed by several numerical simulations, but some results are at odds with the theoretical predictions. Here, I focus on the most common inconsistency, the observation of a seemingly finite conductivity in systems where theory instead predicts a divergence (see e.g., [10]). I analyse the dependence of the flux on the chain length with the help of an approximate formula proposed while discussing the behavior of nearly integrable systems [17]. The good reproduction of the numerical results let me conclude that heat conductivity eventually diverges. In some cases, the seemingly size-independent conductivity is a strong finite-size effect due to the very long phonon mean-free path (MFP) of the modes which carry the energy.
In Sec. 2, the relevant observables are introduced, together with the mathematical and numerical tools. Sec. 3, deals with nearly integrable systems. There, the leading assumption introduced in Ref. [17] is rephrased to facilitate its application in the following sections. In Sec. 4, the behavior of two models of oscillators interacting via non binding forces, is analysed, showing that they behave similarly to the nearly integrable models discussed in the previous section. In Sec. 5, the linear stochastic model proposed by Basile et al. [18] is considered in the small noise limit. It is shown that that no intermediate finite conductivity is therein found and provide an explanation thereof. Sec. 6 is devoted to the study of a variant of the ding-a-ling model [19], which conserves linear momentum and yet was previously found to exhibit normal conductivity [20]. Contrary to the initial claims, I find a clean evidence of a diverging conductivity, in full agreement with the theoretical rate 1/3. Finally, in Sec. 7, some conclusions are drawn and open problems briefly mentioned.
2 Methods and tools
The typical model considered in this paper is ruled by the Hamiltonian
$H=\sum_{n=1}^{L}\left[\frac{p_{n}^{2}}{2m_{n}}+V(q_{n+1}-q_{n})\right]\,$, (1)
where $m_{n}$, $q_{n}$ and $p_{n}$ represent respectively the mass, displacement and momentum of the $n$th particle. $L$ is used both to denote the number of particles as well as the system size (an average unit interparticle distance is implicitly assumed). Everywhere, I work with fixed boundary conditions.
The first (last) particle of the chain is attached to a heat bath operating at temperature $T_{L}$ ($T_{R}$). All of the work is based on the study of the resulting NonEquilibrium Stationary State (NESS) for different chain lengths and, occasionally, different parameter values.
One reason I prefer this approach with respect to the analysis of equilibrium fluctuations is that one can straightforwardly§ identify the appropriate (long) time scale, once a specific system size is given. In fact, given a chain length $L$, it is conceptually easy to determine the required simulation time on the basis of the corresponding statistical fluctuations of the energy flux $J$. A second justification, is that the flux itself can be easily computed from the energy exchanged by the chain with the heat baths, without the need of engaging ourselves with an appropriate bulk definition which can be relatively tricky if the coupling extends beyond nearest neighbours as for the model discussed in Sec. 6.
As for the coupling with the heat baths, it is typically implemented by randomly replacing the velocity of the first and last particle with a zero-average Gaussian-distributed variable whose standard deviation is fixed in accordance to the prescribed temperature. This way, one can use a symplectic algorithm to integrate the equations of motion. The time separation between consecutive thermstat actions quantifies the interaction strength. As already discussed in [2], weak and too strong interactions reduce the flux, making the numerical estimate less statistically reliable. In this paper, time separation is everywhere chosen to be approximately 1 time unit. Moreover, for numerical convenience, it is always constant (I have verified that this choice does not lead to appreciable differences).
A serious obstacle for an accurate estimate of the average flux in large chains is the selection of the initial condition. If it is not sampled according to the (unkwnon) stationary distribution, one must discard a transient to allow the distribution to converge to its asymptotic stationary shape. The larger is $L$ and the longer is this time. Here, I adopt a trick which allows shortening this time as much as possible. The overall analysis of a given model is performed by progressively doubling the chain length so that one can use the outcome of the simulations for size $L$, to generate a proper initial condition for the size $2L$. More precisely, let $\{q_{i}^{(a)},p_{i}^{(a)}\}_{L}$ and $\{q_{i}^{(b)},p_{i}^{(b)}\}_{L}$ be two NESS configurations generated while studying a chain of length $L$ well separated in time to be considered statistically independent. The initial condition $\{\overline{q}_{i},\overline{p}_{i}\}_{2L}$ of the chain of length $2L$ is built by alternating the momenta of configuration $(a)$ with those of configuration $(b)$. This way, the resulting temperature profile is as close as possible to the expected asymptotic one. As for the positions, the procedure is a bit more involved, since the potential energy depends on the relative distances. In practice, I first determine $u_{i}^{(a)}=q_{i}^{(a)}-q_{i-1}^{(a)}$ and $u_{i}^{(b)}=q_{i}^{(b)}-q_{i-1}^{(b)}$ to then obtain $\overline{u}_{i}$ by alternating the two $u_{i}$ sequences and finally generate $\overline{q}_{i}$ by integrating in space the $\overline{u}_{i}$ separations. A final micro adjustment is made to the $\overline{q}_{i}$ variables to ensure the desired chain length. This way, the initial configuration is a good approximation of a generic stationary state and, in fact, simulations do not reveal any appreciable transient in the simulation of the longer chain.
Finally, by denoting with $J(L)$ the heat flux in a system of size $L$ and with $\Delta T=T_{L}-T_{R}$ the temperature difference ($T_{L}$ is assumed to be larger than $T_{R}$), the effective conductivity can be defined as
$\kappa(L)=\frac{J(L)L}{\Delta T}$ (2)
since $\Delta T/L$ is, in fact the temperature gradient imposed by the thermostats. A system is said to satify Fourier law if $\kappa(L)$ is independent of $L$ for $L\to\infty$. If, instead, $\kappa\approx L^{\alpha}$, the system is said to exhibit anomalous conductivity. As anticipated in the introduction, in the widest class of anomalous systems $\alpha=1/3$.
3 Pseudo-normal conductivity
Among the claims of normal conductivity published in the literature, the most surprising are perhaps those which concern nearly integrable systems [21, 22, 23]. The puzzle was solved in Ref. [17], by arguing that in such systems, for not too large chain lengths, the heat resistance can be approximately expressed as the sum of two contributions
$R=R_{0}\left(1+\frac{L}{\it l}\right)$ (3)
where $R_{0}$ can be interpreted as a contact resistance, while $\it l$ is the MFP of the phonons (or, more in general, the modes which transport energy). Roughly speaking, for $L\ll{\it l}$ the behavior is ballistic, i.e. the resistance and thus the flux is independent of $L$. For $L\gg{\it l}$, the system operates in the diffusive regime, characterized by a constant conductivity.
Nevertheless, upon increasing $L$, at some point, nonlinear hydrodynamic mechanisms become relevant, opening a second channel for heat transport. One can include this further process by expressing the heat conductivity as
$\kappa(L)=\frac{1}{\Delta T}\left[\frac{1/R_{0}}{1/L+1/{\it l}}+\sigma L^{\alpha}\right]$ (4)
where the exponent $\alpha$ identifies the universality class. For increasing $L$, the first term in square brackets saturates at ${\it l}/R_{0}$, so that the anomalous component prevails only when the length becomes larger than the crossover value
$L_{c}=\left(\frac{\it l}{\sigma R_{0}}\right)^{1/\alpha}$ (5)
Close to integrability, one expects the MFP to diverge. I assume ${\it l}\approx\varepsilon^{-\theta}$, where $\varepsilon$ parametrizes the distance from the integrable limit, and $\theta$ identifies a model-dependent growth rate of the MFP. Under this further hypothesis,
$L_{c}\approx\varepsilon^{-\theta/\alpha}$ (6)
which can indeed be very large for small $\varepsilon$. In Ref. [17], this approach has been validated by studying, for instance, the diatomic Hard Point Gas (HPG). The HPG consists of point particles which interact only via elastic collisions. If all particles have the same mass, the system is integrable and the energy flux is independent of the system size. If a small asymmetry $\varepsilon$ is introduced between the particle masses on even and odd sites, it turns out that ${\it l}\varepsilon^{-2}$, i.e. $\theta=2$, so that, recalling that $\alpha=1/3$, one finds $L_{c}\approx\varepsilon^{-6}$. As a result, the anomalous behavior may emerge for very large, possibly numerically unaccessible lengths.
4 Non binding potentials
Besides nearly integrable models, a seemingly finite conductivity has been unexpectedly observed in systems of non-binding potentials. Particularly intriguing are the numerical studies of Lennard-Jones, Morse, and Coulomb potentials [24, 25], because of the physical relevance of the models themselves. However, one should also recall that in other instances of non binding models such as the diatomic HPG [26, 27, 28], or a system with a so-called hard-shoulder potential [29], the expected divergence is instead observed.
In order tho shed some light on the overall scenario, I have decided to investigate a softened version of the HPG (SPG: Soft Point Gas), where the interactions are described by the repulsive potential
$V(u)=\frac{C}{u^{2}}$ (7)
where $u$ is the interparticle distance and $C$ has been set equal to 1/2 for easiness of simulations.
The heat-bath temperatures have been set to $T_{L}=1$, and $T_{R}=0.5$, respectively, while the time step has been fixed to $\delta t=0.002$ in order to reproduce the collision dynamics with sufficient accuracy. Furthermore, in analogy with the HPG, I have considered a diatomic version of the SPG, where the mass values alternate between $(1+\varepsilon)$ and $(1-\varepsilon)$. Simulations have been performed for different lengths and different degrees of homogeneity. The fluxes obtained for $\varepsilon=0$, $\varepsilon=0.1$, and $\varepsilon=0.3$ are reported in Fig. 1.
Figure 1: Effective conductivity of the SPG versus system size for three different $\varepsilon$-values: 0, 0.1, and 0.3 from top to bottom. Symbols refer to direct simulations, while the dashed lines are the results of a fit with the expression (2), where $R_{0}$, $\it l$, and $c$ are free parameters. In the inset, the resistance (essentially the inverse of the flux) is reported in the homogeneous case $\varepsilon=0$.
The asymptotic behavior is not clear: at most, one can conjecture that the conductivity tends to saturate in the homogeneous case (see the topmost curve, which corresponds to $\varepsilon=0$). Actually, the behavior in this case is definitely more transparent if one plots the resistance versus $L$. The inset of Fig. 1, in fact, gives evidence of a normal conductivity (see the linear growth, in accordance with Eq. (3)). The reason for the “failed” identification of this behavior in the main body of the figure is the not-so-small contact resistance, identifiable from the crossing of the line of points with the $R$-axis in the inset.
It is thus tempting to use Eq. (4) to describe the behavior of the effective conductivity for the different degrees of mass inhomogeneity. The dashed lines shown in Fig. 1 correspond to the resulting fits, made by considering $R_{0}$, $\it l$, and $\sigma$ as free parameters and setting a priori $\alpha=1/3$ ($\Delta T=0.5$ is known a priori) The agreement is remarkably good in all cases, confirming that the theoretical expression (4) captures the crossover towards the expected divergence, even when the phenomenon itself is not obviously visible in the range of lengths accessible in the simulations.
The fitted parameters can be then used to investigate the dependence of the transport properties on the mass asymmetry $\varepsilon$. The most instructive observable is the MFP: from Fig. 2 (see the upper black curve), one can see that $\it l$ grows significantly while approaching the homogeneous regime, without nevertheless displaying a divergence as in an integrable limit.
Figure 2: The mean free path $\it l$ (black) and the amplitude $\sigma$ (red) of the nonlinear term in the SPG for different degrees of mass inhomogeneity.
In order to shed further light on the possible closeness of the homogeneous case to an integrable dynamics, I have computed the Lyapunov exponents for $\varepsilon=0$ and three different energy densities. The spectra, reported in Fig. 3 have the typical shape observed in chaotic Hamiltonian system. Moreover, the maximum exponent decreases upon decreasing the energy (and hence the nonlinearity) and its numerical values do not give any hint of vicinity to an integrable limit.
Figure 3: Lyapunov spectra of the SPG fro three different energy densities: 0.5, 1, 2, from bottom to top. Only the first half of the spectrum is displayed
Further interesting information is contained in the dependence of $\sigma$ on $\varepsilon$. In Fig. 2, one can see that $\sigma$ becomes increasingly small, when $\varepsilon\to 0$. This contributes to further increase the crossover length $L_{c}$ as from Eq. (5) ($R_{0}$ remains steadily close to 6 when $\varepsilon\to 0$). This is at variance with the diatomic HPG studied in [17], where it is understood that $\sigma$ does not exhibit any particular (pseudo)-singularity in the integrable limit.
Altogether, we have seen that the homogeneous SPG is qualitatively similar In conclusion, we have seen that although the homogeneous SPG is qualitatively similar to the systems characterized by a non binding potential analysed in [24, 25], it is characterized by a diverging conductivity for sufficiently large length scales. In order to strengthen the conclusion that the anomalous divergence is indeed an ubiquituous features within non-binding potentials, I have also revisited the the specific model thoroughly investigated in [25], where the authors have also developed kinetic arguments to determine the value of the finite conductivity. The model is characterized by a repulsive potential: parabolic on one side and flat on the other,
${V(u)=}\begin{cases}(u-d)^{2}/2&u<d\\ 0&u\geq d\end{cases}$ (8)
I investigated the case $d=1$, because as claimed in [25], the conductivity should be independent of the temperature: Again, the boundary temperatures are $T_{L}=1$ and $T_{R}=0.5$. The results are plotted in Fig. 4, by using the same format as in Ref. [25] (i.e. the vertical scale is linear). Up to $L=8192$, the values agree with those in Fig. 3 of Ref. [25] ($L=8192$ was the maximum value reported therein). Above $L=8192$, though slowly, the effective conductivity keeps increasing. A fit with Eq. (4) produces the red curve, which follows very closely the numerical points.
Figure 4: Effective conductivity versus the chain length $L$. Full dots refer to the outcome of direct simulations. The solid red line is the outcome of a fit with expression (4). The upper horizontal dashed line corresponds to the conductivity determined in [25] with kinetic arguments. The lower horizonatal line corresponds to the conductvity extracted from the above mentioned fit.
Once again, it appears that the anomalous divergence is expected to arise in this model as well. On a quantitative level, I now compare the estimate of the (finite) conductivity presented in Ref. [25] with the diffusion component ${\it l}/R_{0}$, estimated from the fit. The two values correspond to the two horizontal lines reported in Fig. 4. There is a certain disagreement, but in spite of the very different approaches adopted to obtain them, I would conclude that they are relatively consistent. In fact, one should note that the numerically estimated diffusive conductivity is based on the assumption that Eq. (4) captures exactly the “short”-length beahvior, but it is well possible that the formula (4) is too simple to estimate the flux over such scales.
5 A stochastic linear system
The stochastic model introduced in Ref. [18] is an interesting testbed for the analysis of regimes close to an integrable limit, since its linear structure allows to perform some analytical calculations.
In its simplest formulation [14], the model consists of a harmonic chain, where the velocities of two randomly chosen consecutive particles are exchanged with a given rate $\varepsilon$. The random update mimics the effect of a nonlinearity by inducing scattering processes among the otherwise uncoupled Fourier modes. In the limit $\varepsilon\to 0$, the Fourier modes travel undisturbed; hence $\varepsilon$ can be seen as parameter controlling the distance from integrability. In this model, the conductivity is known to diverge as $L^{1/2}$ [18]. An exact expression for the asymptotic value of the flux was determined in Ref. [30]
$J=\frac{\Delta T}{8(\sqrt{8}-1)\zeta(3/2)}\sqrt{\frac{\pi^{3}}{\varepsilon L}}=\frac{0.14572\Delta T}{\sqrt{\varepsilon L}}\,.$ (9)
We expect this formula to be correct in the limit $L\to\infty$, but what happens for generic values of $L$? The results obtained for two set of simulations performed for different values of $L$ are reported in Fig. 5, where we see that the dependence on the combined variable $\varepsilon L$ displayed by the asymptotic formula (9) is valid in full generality.
Figure 5: Flux of the stochastic model versus the combined variable $\varepsilon L$ for two different values of $L$: 800 (black dots) and 1600 (red diamonds). $T_{R}=5.5$, $T_{L}=4.5$. The dashed curve corresponds to $0.14572/\sqrt{1+\varepsilon L}$
Quite interestingly, the numerical data are well reproduced by the function $0.1475/\sqrt{1+\varepsilon L}$, equal to the asymptotic expression, the only difference being a shift of the argument by 1 unit. We can conclude that, at variance with all previous cases, here there cannot be a crossover from a diffusive to an anomalous regime and this can be justified in the following way. The end of the ballistic regime is signalled by $\varepsilon L\approx 1$, which means that the MFP is ${\it l}\approx\varepsilon^{-1}$ (i.e. $\theta=1$). On the other hand, the asymptotic formula itself suggests that $\sigma\approx\varepsilon^{-1/2}$ By inserting this scaling Ansatz into the formula (5), we conclude that the hypothetical crossover length $L_{c}\approx\varepsilon^{-1}$ is of the same order as the MFP itself. In other words there canot be an intermediate region where diffusion prevails and, a posteriori, this is also the reason why the splitting assumed in Eq. (4) does not work.
In conclusion, this stochastic model behaves very differently from the other nearly integrable models. The reason is the divergence of the coefficient $\sigma$ while approaching the integrable limit. Is this the signature of an underlying significantly different mechanism?
6 Ding-a-ling type model
The last system I am going to analyse is a variant of the ding-a-ling model [19]. In the ding-a-ling model, two types of particles alternate: even particles are harmonically attached to an external substrate, while odd particles interact with the even ones via elastic collisions. The original model exhibits a finite conductivity and this is well understood, since linear momentum is not conserved. In Ref. [20], a variant was proposed, where each even particle harmonically interacts with the previous and following even particle. Being all interactions internal, linear momentum is conserved and one would now expect a divergence of the heat conductivity. Unexpectedly, the authors of [20] found that Fourier law is satisfied.
A few years later, a variant of the the same model, without hard collisions and a different coupling scheme was proposed and studied, finding evidence of a diverging conductivity [31]. At a first glance, the only possible justification for the strange behavior observed in Ref. [20] is that, at variance with all other simple models, it includes next-to-nearest neighbour interactions, but this does not seem to be a convincing explanation, since the asymptotic behavior is the natural result of a coarse-graining process.
Hence, I have decided to revisit the model. Most of the simulations in Ref. [20] have been performed in an equilibrium setup (by studying fluctuations), and the flux is estimated by referring to a “non-standard” macroscopic formula in terms of Fourier modes. Details of the simulations are scarce, especially about the realization of the NESS. As discussed in Sec. 2, here I focus exclusively on nonequilibrium stationary simulations for various reasons, including the possibility to determine the energy flux in a straightforward way.
Additional details: the chain length is fixed so that the averege equilibrium interparticle distance is 1. The number $L$ of particles is odd: particle 1 ($L$) may collide with a hard wall in 0 ($L+1$) in which case its velocity is randomly reset according to the temperature $T_{L}$ ($T_{R}$) (the velocity distribution is $P(v)\propto v\exp[-v^{2}/(2T)]$ to ensure the right sign). The first and $L$th particle are also attached to the 2nd and (N1)(N-1)st particle respectively, via harmonic springs. Finally, because of the more complex spatial structure, in this case I have not implemented the trick discussed in Sec. 2 to generate optimal initial conditions of length $2L$, starting from final conditions obtained for length $L$: I limited myself to discard a sufficiently long transient.
Simulations have been performed by implementing a simple but approximate scheme: equations are integrated with a symplectic algorithm (bilateral leap-frog). At the end of the each time step, pairs of particles which have crossed each other (because of the finite time step) are rearranged: (i) their velocities are exchanged (to simulate the collision); the position of the free particle is shifted to that of the bound particle, to restore the spatial order. This way, the total linear momentum and the total energy are perfectly conserved (as well as the total stretch). Most of the simulations have been performed by setting the time step equal to $\delta t=10^{-2}$, but some have been repeated for $\delta t=10^{-3}$, confirming the numerical results.
The outcome is shown in Fig. 6, where the temperature profile is presented for a numerical experiment where the heat baths operate at the temperatures $T_{L}=1$ and $T_{R}=0.5$, respectively (see the black curve) The profile is typical of those emerging in the presence of anomalous (diverging) conductivity.
Figure 6: Temperature profile for the hard (black) and soft (red) version of the model in [20] versus the rescaled length $x/L$ (the chain length is 1025).
This expectation is confirmed by the data reported in Fig. 7, where the flux $J$ (black dots) decreases approximately as $L^{-2/3}$ (see the dashed line). This is more quantitatively confirmed in the inset, where we see that $JL^{2/3}$ saturates and remains constant. The exponent $2/3$ is the value which corresponds to $\alpha=1/3$: the characteristic rate of the universality class of asymmetric effective potentials.
Since these results contradict the claims made in [20], I have also investigated a softened version of the model, where the interactions between free and harmonic particles are ruled by the smooth repulsive potential (7), where $C$ has been set equal $10^{-3}$ to make it resemble hard elastic collisions. The advantage of this model is that there is no need to accurately identify the collision times; the disadvantage is that the integration time step must be very small, in order to describe accurately the bouncing process (the time step was $\delta t=5\ 10^{-4}$).
The outcome of the simulations is reported again in Figs. 6,7. The temperature profile reported in the first figure (red curve) is slightly lower than in the previous case, but the overall shape is very similar. As for the energy flux reported in the second figure (red stars), we even observe a semi-quantitative agreement. The most important point is, however, that heat conductivity behaves again in an anomalous way, in agreement with the theoretical expectations. The motivation for the different results reported in [20] is unclear. Nevertheless the inconsistency between the behavior of this ding-dong type model and the theoretical expectations now disappears.
Figure 7: Energy flux versus the chain length for the soft and the hard version of the ding-a-ling type model of Ref. [20] (in the inset the flux is scaled by a factor $L^{2/3}$).
7 Conclusion
Fluctuating hydrodynamics shows that three conservation laws (energy, linear momentum, and stretch) are a prerequisite for the emergence of anomalous heat conductivity in one dimension. Several publications have anyhow challenged this scenario, providing examples of a finite conductivity in models where a divergence should be instead observed. I have revisited some of them, showing that these previous conclusions are based on finite-size effects: (much) longer system-sizes are needed in order to see the divergence predicted by theory. The observation of a finite conductivity in the model of Ref. [20] is an exception, presumably due to an error in the estimate of the total flux.
Now, that several inconsistencies between predictions and observartions have been removed, a question still stands about the emergence, in some models, of a seemingly finite conductivity over a broad range of scales. The answer is straightforward while dealing with nearly integrable systems, but it is much less so when the model itsels is manifestly chaotic, as it happens for the soft point chain analysed in this paper. Moreover, not all non-binding potentials exhibit a nearly constant effective conductivity.
An additional complication comes from the amplitude of term gauging the strength of the diverging contribution; it is very small in the homogeneous soft point chain, while it diverges in the stochastic linear model when it approaches the integrable limit. How to justify such a wide range of different results?
The author thanks the StatPhys29 organizers, since their invitation has obliged him to revisit many awkward results which appeared in the last years in the literature.

### v2

@@
 Abstract
-Numerical studies of some unidimensional systems provide evidence of finite thermal conductivity, where theory predicts a divergence with the system size.
-Some models are here reviewed under the working hypothesis that the energy flux across a nonequilibrium stationary state may be the sum of two contributions: the former responsible for normal conductivity, the latter accounting for the anomalous component.
-I conclude that in the thermodynamic limit, the diverging component is indeed present, although the crossover to a regime dominated by the anomalous component may occur at extremely long system sizes.
-Finally, I study a variant of the ding-a-ling model, previously claimed to satisfy Fourier law, showing that it too exhibits a diverging conductivity in the thermodynamic limit.
+Numerical studies of some unidimensional systems suggest that Fourier law is satisfied, where theory predicts a divergence of heat conductivity with the system size.
+Here, I revisit some such models, finding that in all cases a divergence asymptotically emerges.
+This includes a variant of the ding-a-ling model, where I find that, contrary to previous claims, the “anomalous” growth starts already for moderate system sizes.
+More conceptually interesting is the case of non-binding potentials, whose behavior is well reproduced by assuming that the energy flux across the nonequilibrium stationary state is the sum of two contributions: a diffusive and a hydrodynamic one.
+This approach, which extends an idea previously formulated for nearly integrable systems, allows to conclude that the asymptotic regime is always dominated by the anomalous hydrodynamic component, but the crossover may occur for extremely long system sizes.
 1 Introduction
@@
 The general scenario has been confirmed by several numerical simulations, but some results are at odds with the theoretical predictions.
-Here, I focus on the most common inconsistency, the observation of a seemingly finite conductivity in systems where theory instead predicts a divergence (see e.g., [10]).
-I analyse the dependence of the flux on the chain length with the help of an approximate formula proposed while discussing the behavior of nearly integrable systems [17].
-The good reproduction of the numerical results let me conclude that heat conductivity eventually diverges.
-In some cases, the seemingly size-independent conductivity is a strong finite-size effect due to the very long phonon mean-free path (MFP) of the modes which carry the energy.
+Here, I focus on the typical inconsistency: the evidence of a seemingly finite conductivity in cases where theory predicts a divergence (see e.g., [10]).
+I show that many numerical observations are affected by strong finite-size effects, which mask the asymptotic divergence.
+In practice, in many models, the effective heat conductivity is the sum of two contributions, a kinetic one sustained by phonon-like modes characterized by a long mean-free path (MFP).
+This idea was already proposed in [17] to explain heat transport in nearly integrable systems.
+Therein, it is natural to expect phonon-like modes to propagate over long distances before dissipating their energy.
+Here, I show that the same decomposition applies to a much wider class of systems, namely models characterized by non-binding energy which are not nearly integrable, as revealed by their fully chaotic spectrum of Lyapunov exponents.
 In Sec.
 2, the relevant observables are introduced, together with the mathematical and numerical tools.
+In Sec.
+3, I briefly revisit the conjecture introduced in Ref.
+[17] to describe the behavior of nearly integrable systems, rephrasing it, to introduce its application in the following section.
+In Sec.
+4, the behavior of two models of oscillators interacting via non binding forces, is analysed, showing that the effective conductivity is, to a large extent, the sum of a normal and an diverging component.
 Sec.
-3, deals with nearly integrable systems.
-There, the leading assumption introduced in Ref.
-[17] is rephrased to facilitate its application in the following sections.
-In Sec.
-4, the behavior of two models of oscillators interacting via non binding forces, is analysed, showing that they behave similarly to the nearly integrable models discussed in the previous section.
-In Sec.
-5, the linear stochastic model proposed by Basile et al.
-[18] is considered in the small noise limit.
-It is shown that that no intermediate finite conductivity is therein found and provide an explanation thereof.
-Sec.
-6 is devoted to the study of a variant of the ding-a-ling model [19], which conserves linear momentum and yet was previously found to exhibit normal conductivity [20].
-Contrary to the initial claims, I find a clean evidence of a diverging conductivity, in full agreement with the theoretical rate 1/3.
+5 is devoted to the study of a variant of the ding-a-ling model [18], which conserves linear momentum and is often quoted as a clear example of normal conductivity [19].
+Contrary to the initial claims, I find compelling evidence of a diverging conductivity, in full agreement with the theoretical expectation $L^{1/3}$.
 Finally, in Sec.
-7, some conclusions are drawn and open problems briefly mentioned.
+6, some conclusions are drawn and two open problems briefly mentioned.
 2 Methods and tools
@@
 A second justification, is that the flux itself can be easily computed from the energy exchanged by the chain with the heat baths, without the need of engaging ourselves with an appropriate bulk definition which can be relatively tricky if the coupling extends beyond nearest neighbours as for the model discussed in Sec.
-6.
+5.
 As for the coupling with the heat baths, it is typically implemented by randomly replacing the velocity of the first and last particle with a zero-average Gaussian-distributed variable whose standard deviation is fixed in accordance to the prescribed temperature.
@@
 3 Pseudo-normal conductivity
-Among the claims of normal conductivity published in the literature, the most surprising are perhaps those which concern nearly integrable systems [21, 22, 23].
-The puzzle was solved in Ref.
-[17], by arguing that in such systems, for not too large chain lengths, the heat resistance can be approximately expressed as the sum of two contributions
-$R=R_{0}\left(1+\frac{L}{\it l}\right)$ (3)
-where $R_{0}$ can be interpreted as a contact resistance, while $\it l$ is the MFP of the phonons (or, more in general, the modes which transport energy).
-Roughly speaking, for $L\ll{\it l}$ the behavior is ballistic, i.e.
-the resistance and thus the flux is independent of $L$.
-For $L\gg{\it l}$, the system operates in the diffusive regime, characterized by a constant conductivity.
-Nevertheless, upon increasing $L$, at some point, nonlinear hydrodynamic mechanisms become relevant, opening a second channel for heat transport.
-One can include this further process by expressing the heat conductivity as
-$\kappa(L)=\frac{1}{\Delta T}\left[\frac{1/R_{0}}{1/L+1/{\it l}}+\sigma L^{\alpha}\right]$ (4)
-where the exponent $\alpha$ identifies the universality class.
-For increasing $L$, the first term in square brackets saturates at ${\it l}/R_{0}$, so that the anomalous component prevails only when the length becomes larger than the crossover value
+Nearly integrable systems represent a class of models which, surprisingly, seem to exhibit normal conductivity [20, 21, 22].
+An explanation of this anomaly was given in Ref.
+[17], by arguing that in such systems, the heat resistance can be approximately expressed as the sum of two contributions: (i) a kinetic term arising from the propagation of phonon-like modes; (ii) a hydrodynamic term, responsible for the asymptotic divergence of the conductivity.
+Here, I review the method, underlining the relevant conceptual assumptions.
+The simplest description of the kinetic channel is, in terms of the effective resistance,
+$R_{kin}=R_{0}\left(1+\frac{L}{\it l}\right)\;.$ (3)
+Here $R_{0}$ can be interpreted as a contact resistance, due to the impedence mismatch between the thermal bath and the physical system (see also [23]); $\it l$ is, instead, the MFP of the phonons.
+For $L\ll{\it l}$, the behavior is ballistic, i.e.
+the resistance and hence the flux is independent of $L$.
+For $L\gg{\it l}$, the system operates in the diffusive regime, characterized by a resistance proportional to the system length.
+The anomalous component, whenever present, consists of a resistance which grows sublinearly with $L$, i.e.
+$R_{an}=\frac{L^{1-\alpha}}{\sigma}\;,$ (4)
+where $\alpha>0$ identifies the universality class, while $\sigma$ gauges the overall amplitude of this term.
+The core of the conjecture is that for finite lengths, the two channels of heat transport operate in parallel, so that the effective conductivity $\kappa(L)$ is the sum of the two contributions.
+More precisely, in the limit of small temperature differences,
+$\kappa(L)=\frac{1}{(\Delta T)L}\left[\frac{1}{R_{kin}}+\frac{1}{R_{an}}\right]=\frac{1}{\Delta T}\left[\frac{1/R_{0}}{1/L+1/{\it l}}+\sigma L^{\alpha}\right]$ (5)
+The first term in square brackets saturates at ${\it l}/R_{0}$, so that the anomalous component prevails only when the length is larger than the crossover value
 $L_{c}=\left(\frac{\it l}{\sigma R_{0}}\right)^{1/\alpha}$ (6)
-Close to integrability, one expects the MFP to diverge.
-I assume ${\it l}\approx\varepsilon^{-\theta}$, where $\varepsilon$ parametrizes the distance from the integrable limit, and $\theta$ identifies a model-dependent growth rate of the MFP.
-Under this further hypothesis,
-$L_{c}\approx\varepsilon^{-\theta/\alpha}$ (6)
-which can indeed be very large for small $\varepsilon$.
-In Ref.
-[17], this approach has been validated by studying, for instance, the diatomic Hard Point Gas (HPG).
-The HPG consists of point particles which interact only via elastic collisions.
-If all particles have the same mass, the system is integrable and the energy flux is independent of the system size.
-If a small asymmetry $\varepsilon$ is introduced between the particle masses on even and odd sites, it turns out that ${\it l}\varepsilon^{-2}$, i.e.
-$\theta=2$, so that, recalling that $\alpha=1/3$, one finds $L_{c}\approx\varepsilon^{-6}$.
-As a result, the anomalous behavior may emerge for very large, possibly numerically unaccessible lengths.
+Depending on the various parameter values, $L_{c}$ may be so large to make the detection of the asymptotic divergence practically impossible.
+Nearly integrable systems represent a prominent instance where this assumption has been validated, since therein the MFP $\it l$ can be tuned by adjusting the closeness to integrability.
+More precisely, quantitative studies have been performed in Ref.
+[17], by analysing the diatomic Hard Point Gas (HPG), a system of point particles, which interact only via elastic collisions.
+In that model, when the masses are all equal, the dynamics is integrable and energy is ballistically transported across the chain; a small mass heterogeneity induces a weak energy “dissipation”, and hence a finite MFP.
 4 Non binding potentials
-Besides nearly integrable models, a seemingly finite conductivity has been unexpectedly observed in systems of non-binding potentials.
-Particularly intriguing are the numerical studies of Lennard-Jones, Morse, and Coulomb potentials [24, 25], because of the physical relevance of the models themselves.
-However, one should also recall that in other instances of non binding models such as the diatomic HPG [26, 27, 28], or a system with a so-called hard-shoulder potential [29], the expected divergence is instead observed.
-In order tho shed some light on the overall scenario, I have decided to investigate a softened version of the HPG (SPG: Soft Point Gas), where the interactions are described by the repulsive potential
+In this section, I show that the approach described above applies to a wider class of dynamical systems, far from the integrable limit as testified by their clearly chaotic spectrum of Lyapunov exponents.
+I refer to systems characterized by a non-binding potentials.
+Particularly intriguing are the Lennard-Jones, Morse, and Coulomb potentials [24, 25], because of the physical relevance of these models.
+I first investigate a softened version of the HPG (SPG: Soft Point Gas), characterized by the repulsive potential
 $V(u)=\frac{C}{u^{2}}$ (7)
 where $u$ is the interparticle distance and $C$ has been set equal to 1/2 for easiness of simulations.
-The heat-bath temperatures have been set to $T_{L}=1$, and $T_{R}=0.5$, respectively, while the time step has been fixed to $\delta t=0.002$ in order to reproduce the collision dynamics with sufficient accuracy.
-Furthermore, in analogy with the HPG, I have considered a diatomic version of the SPG, where the mass values alternate between $(1+\varepsilon)$ and $(1-\varepsilon)$.
-Simulations have been performed for different lengths and different degrees of homogeneity.
-The fluxes obtained for $\varepsilon=0$, $\varepsilon=0.1$, and $\varepsilon=0.3$ are reported in Fig.
-1.
+The time step has been fixed to $\delta t=0.002$ in order to reproduce the collision dynamics with sufficient accuracy.
+At variance with the HPG, a homogeneous chain is chaotic (hence non integrable), as testified by the spectrum of Lyapunov exponents: see Fig.
+1, where the first half of the spectrum is displayed for three different energy densities.
+The Lyapunov exponents are clearly larger than 0 and the overall shape is akin to that of standard chaotic Hamiltonian systems.
+Figure 1: Lyapunov spectra of the SPG for three different energy densities: 0.5, 1, 2, from bottom to top. The chain length is $L=100$ and only the first half of the spectrum is displayed
+In spite of the system being chaotic, simulations performed by setting the heat-bath temperatures at $T_{L}=1$, and $T_{R}=0.5$, suggest evidence of a normal conduction, analogously to what observed in Refs.
+[24, 25].
+In fact, the effective conductivity displayed in Fig.
+2 (see the triangles) suggests a slow saturation.
+The evidence is even more compelling in the inset, where the actual resistance, plotted for various system sizes, displays a clean linear growth.
+Hence, I have decided to introduce a small asimmetry between the mass of even ($1+\varepsilon$) and odd ($1-\varepsilon$) particles.
+In fact, as seen in [17] with reference to the HPG, a small $\varepsilon$ is able to downgrade the ballistic transport, allowing for an easier detection of the diverging hydrodynamic component.
 Figure 2: Effective conductivity of the SPG versus system size for three different $\varepsilon$-values: 0, 0.1, and 0.3 from top to bottom.
-Symbols refer to direct simulations, while the dashed lines are the results of a fit with the expression (2), where $R_{0}$, $\it l$, and $c$ are free parameters.
+Symbols refer to direct simulations, while the dashed lines are the results of a fit with the expression (2), where $R_{0}$, $\it l$, and $\sigma$ are free parameters (see Fig.
+3 for the values of $\it l$ and $\sigma$; $R_{0}\approx 6$ in the whole range of $\varepsilon$ values).
 In the inset, the resistance (essentially the inverse of the flux) is reported in the homogeneous case $\varepsilon=0$.
-The asymptotic behavior is not clear: at most, one can conjecture that the conductivity tends to saturate in the homogeneous case (see the topmost curve, which corresponds to $\varepsilon=0$).
-Actually, the behavior in this case is definitely more transparent if one plots the resistance versus $L$.
-The inset of Fig.
-1, in fact, gives evidence of a normal conductivity (see the linear growth, in accordance with Eq.
-(3)).
-The reason for the “failed” identification of this behavior in the main body of the figure is the not-so-small contact resistance, identifiable from the crossing of the line of points with the $R$-axis in the inset.
+The resulting effective conductivities for $\varepsilon=0.1$, and $\varepsilon=0.3$ are again reported in Fig.
+2 (see dots and squares, respectively).
+One can see more evidence of a grwoth, but the asymptotic behavior is not clear.
 It is thus tempting to use Eq.
-(4) to describe the behavior of the effective conductivity for the different degrees of mass inhomogeneity.
+(5) to characterize the effective conductivity for the different degrees of mass inhomogeneity.
 The dashed lines shown in Fig.
@@
 Figure 3: The mean free path $\it l$ (black) and the amplitude $\sigma$ (red) of the nonlinear term in the SPG for different degrees of mass inhomogeneity.
-In order to shed further light on the possible closeness of the homogeneous case to an integrable dynamics, I have computed the Lyapunov exponents for $\varepsilon=0$ and three different energy densities.
-The spectra, reported in Fig.
-3 have the typical shape observed in chaotic Hamiltonian system.
-Moreover, the maximum exponent decreases upon decreasing the energy (and hence the nonlinearity) and its numerical values do not give any hint of vicinity to an integrable limit.
-Figure 3: Lyapunov spectra of the SPG fro three different energy densities: 0.5, 1, 2, from bottom to top. Only the first half of the spectrum is displayed
 Further interesting information is contained in the dependence of $\sigma$ on $\varepsilon$.
 In Fig.
-2, one can see that $\sigma$ becomes increasingly small, when $\varepsilon\to 0$.
+3, one can see that $\sigma$ becomes smaller and smaller , when $\varepsilon\to 0$.
 This contributes to further increase the crossover length $L_{c}$ as from Eq.
@@
 This is at variance with the diatomic HPG studied in [17], where it is understood that $\sigma$ does not exhibit any particular (pseudo)-singularity in the integrable limit.
-Altogether, we have seen that the homogeneous SPG is qualitatively similar In conclusion, we have seen that although the homogeneous SPG is qualitatively similar to the systems characterized by a non binding potential analysed in [24, 25], it is characterized by a diverging conductivity for sufficiently large length scales.
-In order to strengthen the conclusion that the anomalous divergence is indeed an ubiquituous features within non-binding potentials, I have also revisited the the specific model thoroughly investigated in [25], where the authors have also developed kinetic arguments to determine the value of the finite conductivity.
+In conclusion, one can conclude that the homogeneous SPG ($\varepsilon=0)$ is characterized by a diverging conductivity, but a direct measure is prohibitive, since the crossover length (i.e.
+the length where the anomalous component starts being comparable with the kinetic contribution) is about $2\ 10^{4}$.
+In order to strengthen the conclusion that the anomalous divergence is an ubiquituous feature within non-binding potentials, I have also revisited the model thoroughly investigated in [25], where the authors have also developed kinetic arguments to estimate the presumed finite conductivity.
 The model is characterized by a repulsive potential: parabolic on one side and flat on the other,
@@
 In fact, one should note that the numerically estimated diffusive conductivity is based on the assumption that Eq.
-(4) captures exactly the “short”-length beahvior, but it is well possible that the formula (4) is too simple to estimate the flux over such scales.
-5 A stochastic linear system
-The stochastic model introduced in Ref.
-[18] is an interesting testbed for the analysis of regimes close to an integrable limit, since its linear structure allows to perform some analytical calculations.
-In its simplest formulation [14], the model consists of a harmonic chain, where the velocities of two randomly chosen consecutive particles are exchanged with a given rate $\varepsilon$.
-The random update mimics the effect of a nonlinearity by inducing scattering processes among the otherwise uncoupled Fourier modes.
-In the limit $\varepsilon\to 0$, the Fourier modes travel undisturbed; hence $\varepsilon$ can be seen as parameter controlling the distance from integrability.
-In this model, the conductivity is known to diverge as $L^{1/2}$ [18].
-An exact expression for the asymptotic value of the flux was determined in Ref.
-[30]
-$J=\frac{\Delta T}{8(\sqrt{8}-1)\zeta(3/2)}\sqrt{\frac{\pi^{3}}{\varepsilon L}}=\frac{0.14572\Delta T}{\sqrt{\varepsilon L}}\,.$ (9)
-We expect this formula to be correct in the limit $L\to\infty$, but what happens for generic values of $L$?
-The results obtained for two set of simulations performed for different values of $L$ are reported in Fig.
-5, where we see that the dependence on the combined variable $\varepsilon L$ displayed by the asymptotic formula (9) is valid in full generality.
-Figure 5: Flux of the stochastic model versus the combined variable $\varepsilon L$ for two different values of $L$: 800 (black dots) and 1600 (red diamonds).
-$T_{R}=5.5$, $T_{L}=4.5$.
-The dashed curve corresponds to $0.14572/\sqrt{1+\varepsilon L}$
-Quite interestingly, the numerical data are well reproduced by the function $0.1475/\sqrt{1+\varepsilon L}$, equal to the asymptotic expression, the only difference being a shift of the argument by 1 unit.
-We can conclude that, at variance with all previous cases, here there cannot be a crossover from a diffusive to an anomalous regime and this can be justified in the following way.
-The end of the ballistic regime is signalled by $\varepsilon L\approx 1$, which means that the MFP is ${\it l}\approx\varepsilon^{-1}$ (i.e.
-$\theta=1$).
-On the other hand, the asymptotic formula itself suggests that $\sigma\approx\varepsilon^{-1/2}$ By inserting this scaling Ansatz into the formula (5), we conclude that the hypothetical crossover length $L_{c}\approx\varepsilon^{-1}$ is of the same order as the MFP itself.
-In other words there canot be an intermediate region where diffusion prevails and, a posteriori, this is also the reason why the splitting assumed in Eq.
-(4) does not work.
-In conclusion, this stochastic model behaves very differently from the other nearly integrable models.
-The reason is the divergence of the coefficient $\sigma$ while approaching the integrable limit.
-Is this the signature of an underlying significantly different mechanism?
+(5) captures exactly the “short”-length beahvior, but it is well possible that the formula (5) is too simple to reproduce perfectly well the flux over such scales.
 5 Ding-a-ling type model
-The last system I am going to analyse is a variant of the ding-a-ling model [19].
+Besides the class of models with non binding potentials, there exists an awkward result on a variant of the ding-a-ling model [18].
 In the ding-a-ling model, two types of particles alternate: even particles are harmonically attached to an external substrate, while odd particles interact with the even ones via elastic collisions.
@@
 [19] is that, at variance with all other simple models, it includes next-to-nearest neighbour interactions, but this does not seem to be a convincing explanation, since the asymptotic behavior is the natural result of a coarse-graining process.
-Hence, I have decided to revisit the model.
+I have decided to revisit the model, since it is considered a clear counterexample to the theoretical arguments.
 Most of the simulations in Ref.
-[20] have been performed in an equilibrium setup (by studying fluctuations), and the flux is estimated by referring to a “non-standard” macroscopic formula in terms of Fourier modes.
+[19] have been performed in an equilibrium setup (by studying fluctuations), and the flux is estimated by implementing a “non-standard” macroscopic formula expressed in terms of Fourier modes.
 Details of the simulations are scarce, especially about the realization of the NESS.
 As discussed in Sec.
-2, here I focus exclusively on nonequilibrium stationary simulations for various reasons, including the possibility to determine the energy flux in a straightforward way.
-Additional details: the chain length is fixed so that the averege equilibrium interparticle distance is 1.
+2, here I focus exclusively on nonequilibrium stationary simulations for various reasons, including the opportunity to determine the energy flux in a straightforward way.
+Additional details: the chain length is fixed so that the average equilibrium interparticle distance is 1.
 The number $L$ of particles is odd: particle 1 ($L$) may collide with a hard wall in 0 ($L+1$) in which case its velocity is randomly reset according to the temperature $T_{L}$ ($T_{R}$) (the velocity distribution is $P(v)\propto v\exp[-v^{2}/(2T)]$ to ensure the right sign).
@@
 The outcome is shown in Fig.
-6, where the temperature profile is presented for a numerical experiment where the heat baths operate at the temperatures $T_{L}=1$ and $T_{R}=0.5$, respectively (see the black curve) The profile is typical of those emerging in the presence of anomalous (diverging) conductivity.
+5, where the temperature profile is presented for a numerical experiment where the heat baths operate at the temperatures $T_{L}=1$ and $T_{R}=0.5$, respectively (see the black curve) The profile is typical of those emerging in the presence of anomalous (diverging) conductivity.
 Figure 5: Temperature profile for the hard (black) and soft (red) version of the model in [19] versus the rescaled length $x/L$ (the chain length is 1025).
 This expectation is confirmed by the data reported in Fig.
-7, where the flux $J$ (black dots) decreases approximately as $L^{-2/3}$ (see the dashed line).
+6, where the flux $J$ (black dots) decreases approximately as $L^{-2/3}$ (see the dashed line).
 This is more quantitatively confirmed in the inset, where we see that $JL^{2/3}$ saturates and remains constant.
@@
 The outcome of the simulations is reported again in Figs.
-6,7.
+5,6.
 The temperature profile reported in the first figure (red curve) is slightly lower than in the previous case, but the overall shape is very similar.
@@
 The most important point is, however, that heat conductivity behaves again in an anomalous way, in agreement with the theoretical expectations.
-The motivation for the different results reported in [20] is unclear.
-Nevertheless the inconsistency between the behavior of this ding-dong type model and the theoretical expectations now disappears.
+The motivation for the different results reported in [19] is unclear; possibly an improper definition of the local heat flux.
+Nevertheless, it is now clear that this ding-a-ling type model exhibits an anomalous, diverging conductivity.
 Figure 6: Energy flux versus the chain length for the soft and the hard version of the ding-a-ling type model of Ref. [19] (in the inset the flux is scaled by a factor $L^{2/3}$).
@@
 Fluctuating hydrodynamics shows that three conservation laws (energy, linear momentum, and stretch) are a prerequisite for the emergence of anomalous heat conductivity in one dimension.
-Several publications have anyhow challenged this scenario, providing examples of a finite conductivity in models where a divergence should be instead observed.
-I have revisited some of them, showing that these previous conclusions are based on finite-size effects: (much) longer system-sizes are needed in order to see the divergence predicted by theory.
-The observation of a finite conductivity in the model of Ref.
-[20] is an exception, presumably due to an error in the estimate of the total flux.
-Now, that several inconsistencies between predictions and observartions have been removed, a question still stands about the emergence, in some models, of a seemingly finite conductivity over a broad range of scales.
-The answer is straightforward while dealing with nearly integrable systems, but it is much less so when the model itsels is manifestly chaotic, as it happens for the soft point chain analysed in this paper.
-Moreover, not all non-binding potentials exhibit a nearly constant effective conductivity.
-An additional complication comes from the amplitude of term gauging the strength of the diverging contribution; it is very small in the homogeneous soft point chain, while it diverges in the stochastic linear model when it approaches the integrable limit.
-How to justify such a wide range of different results?
+Several publications have anyhow challenged this scenario, providing examples of finite conductivity in models where a divergence should be instead observed.
+I have revisited some of them, showing that in all cases heat conductivity diverges in the thermodynamic limit.
+In the case of the ding-a-ling type model studied in Ref.
+[19], the discrepancy is due to some kind of error presumably related to the definition of the local flux.
+More interesting is the case of non-binding potentials, where I find that the conductivity is affected by very strong finite-size effects, which mask the asymptotic behavior.
+This conclusion is supported by the observation that the effective conductivity can be well approximated as the sum of two distinct contributions: a kinetic diffusive term and the anomalous one.
+This assumption, previously proposed to explain the behavior of nearly integrable systems, proves, indeed, more general and is valid even in fully chaotic models,
+As a result, a question opens up: how can it be that some modes can propagate over very long distances in the presence of an evident non-integrability?
+In a sense, this unexpected scenario is complentary to the anomalous (non ballistic) divergence observed in the symmetric gard-point chain, where the integrals of motions (the velocitons) diffuse rather than propagating ballistically.
+An explanation of when and how this scenario emerges in non-binding potential is a point that will be worth exploring.
+Finally, while representing the conductivity as the sum of two terms, I have also found that the amplitude $\sigma$ of the hydrodynamic term plays a crucial role in determining the crossover length.
+In one of the non-binding potentials, it is very small, thus contributing to further increase the crossover length.
+Understanding which dynamical properties are responsible for the amplitude of $\sigma$ is another point that should be clarified.
 The author thanks the StatPhys29 organizers, since their invitation has obliged him to revisit many awkward results which appeared in the last years in the literature.